You plan to retire 30 years from now. You will save $5,000 per year for the next 15 years, and then $7,000 for the 15 years after that. You can earn 9% on your savings. How much will you have 30 years from now? Round your answer tothe nearest whole number,for example 1441.You will save $4,000 per year for the next 30 years. You will earn 10% on your savings. At the end of 30 years, you will retire, and you expect to live for 40 years after that. How much can you withdraw every year for 40 years and not run out of money? Round your answer to the nearest whole number, for example 400
Answer #1
1) The FV of the annuity of $5000 for 15 years at t15 = 5000*(1.09^15-1)/0.09 = $   1,46,805
FV of the above amount at t30 = 146805*1.09^15 = $   5,34,735
FV of the annuity (t16 to t30) at t30 = 7000*(1.09^15-1)/0.09 = $   2,05,526
Amount accumulated in the savings account 30 years from now = $   7,40,261
2) The FV of the annual savings = 4000*(1.1^30-1)/0.10 = $   6,57,976
The above amount will be the PV of the annual withdrawals for the next 40 years.
So, the annual withdrawals that can be made = 657976*0.10*1.1^40/(1.10^40-1) = $       67,284